calculate the energy required to heat 46g of ice from -15 C to 115 C
1
Expert's answer
2020-12-03T14:03:34-0500

Q = heat to heat ice to 0°C + heat to convert ice to water + heat to heat water + heat to convert water to steam + heat to heat steam


Q = mc∆T + mLf + mc∆T + mLv +mc∆T

Q = m(c∆T + Lf + c∆T + Lv + c∆T)

Q = 46/1000((15×4182)+(333000)+(100×4182)+(2260000)+(15×4182))

Q = 0.046(62730 + 333000 + 418200 + 2260000 + 62730)

Q = 0.046(3136660)

Q = 144286.36J

Q = 144.3kJ